# GaSLight: Gaussian Splats for Spatially-Varying Lighting in HDR

### v2

@@ 4.2 Light source estimation
Method MSE PSNR MSE PSNR
LDR $0.0715$ $13.0605$ $0.0282$ $17.5467$
HDRCNN $0.0246$ $23.6138$ $0.0060$ $30.6715$
MaskHDR $0.0243$ $23.5503$ $0.0061$ $30.9510$
SingleHDR $0.0178$ $24.0680$ $0.0092$ $23.6101$
ExpandNet $0.0187$ $24.7158$ $0.0148$ $22.5799$
Exposure Diff $0.0264$ $23.7384$ $0.0061$ $27.4963$
Ours $0.0131$ $25.3895$ $0.0041$ $28.7022$
Table 3: Quantitative evaluation of renderings lit by each method’s HDR image output. Results are color coded by best, second and third best.
We evaluate our novel light intensity estimation method (sec. 3.1) on the task of bright light source prediction. We first show qualitative results on the filtered Si-HDR dataset in fig. 4 and on the BtP-HDR dataset in fig. 5. The figures illustrate both the luminance of the generated HDR maps, as well as a render of a sphere on a plane lit by the color HDR map. Since these images are not full 360° environment maps, we project them onto a partial hemisphere of 90° field of view, and use this as light source for rendering. The camera, plane, and sphere are located at the origin, facing the camera. The glossy sphere has an albedo of 1 (pure white) and a roughness coefficient of 0. We make sure all HDR images used as lighting are exposed the same way by aligning the averages of pixels with intensity between $0.2$ and $0.8$ with the ground truth. Observe how the predictions made by our method more closely match the ground truth, especially when bright light sources generate hard cast shadows on the plane and bright specularities on the sphere.
To further assess the accuracy of our method, we compute quantitative image comparison metrics (MSE and PSNR) on the renders compared to the ground truth in tab. 3. We observe that our method significantly outperforms existing methods in terms of both MSE and PSNR on the filtered Si-HDR dataset. On the BtP-HDR dataset, our method provides the best MSE score while offering a competitive PSNR value. We attribute this lower score to the fact that BtP-HDR features indoor scenes that tend to have a lower dynamic range than outdoor environments. Consequently, techniques like HDRCNN and MaskHDR perform well in areas that are slightly above saturation.
@@
PanoHDR-NeRF Ours
Scene MSE PSNR MSE PSNR
Chess room $0.0247$ $17.5010$ $0.0053$ $25.9313$
Small class $0.1097$ $9.7872$ $0.0061$ $23.1244$
Table 4: Quantitative evaluation of renderings lit by novel views of PanoHDR-NeRF [19] and Ours, as compared to the captured HDR ground truth panoramas provided in [19]. Our method achieves the best scores on both scenes for all metrics.
@@
Input LDR HDRCNN MaskHDR SingleHDR ExpandNet Exposure Diff ours GT
Figure 5: Qualitative evaluation on the BtP-HDR dataset [5]. From an LDR image (left), we compare the luminance of each method’s predicted HDR (even rows) and the rendering of a glossy sphere illuminated by the prediction (odd rows). Our approach more accurately estimates light intensity, producing renders similar to those rendered using the ground truth HDR image.
@@
5 Application: virtual object insertion
The 3DGS representation is inherently 3D, allowing for near-field lighting effects out-of-the-box. It is possible to composite virtual objects directly in this representation, as shown in LABEL:fig:teaser. We obtain these results by importing our 3DGS representation directly in Blender [8], setting the Gaussians as emitters, and rendering the scene using the Cycles physically-based renderer. The inserted objects are automatically lit with near-field lighting effects, without manual intervention or custom lights in the scene.
@@ 6 Conclusion
Despite encouraging results, GaSLight suffers from some limitations. First, each Gaussian in the 3DGS representation is effectively considered to be a light source (e.g., when rendering, see sec. 5. This precludes the simulation of more complex lighting effects, such as interreflections or scattering. Here, it is likely that new raytracing-based approaches [40] could help. Second, our HDR light intensity estimation network is not trained to achieve multi-view consistency: it is therefore possible that different content is extrapolated for two nearby views. While the spherical harmonics in the 3DGS representation alleviates for this somewhat and can realistically blend inconsistent predictions, future work could enforce such consistency, for example through score distillation sampling [46]. We also note that, as opposed to most lighting estimation methods that predict light sources outside the field of view of the image, we only recover the lighting for visible parts of the scene. Combining our method with view extrapolation approaches (e.g., [7, 43]) is a promising future direction. Finally, our current implementation, which relies on InstantSplat [14] does not scale well to large scenes. This is somewhat orthogonal to our goal of capturing lighting—we are hopeful that new developments (e.g., [30]) could help alleviate this.
We present GaSLight, a novel Gaussian Splatting-based HDR Lighting representation. The core of our method is a diffusion-based LDR-to-HDR model that uses its priors of natural images to expand the dynamic ranges of LDR images to unsaturated, even on bright light sources such as the sun. We further fit a 3DGS model on the predicted HDR images, which results in a light representation that can model spatially-varying effects, achieves high frequency reflections and hard cast shadows, and can readily be integrated in existing rendering engines.
Acknowledgements  This research was supported by Sentinel North, NSERC grant RGPIN 2020-04799, and the Digital Research Alliance Canada.

### v3

@@ 4.2 Light source estimation
Relight HDR Relight HDR
Method MSE PSNR Ang err H-VDP3 P-PIQE MSE PSNR Ang err H-VDP3 P-PIQE
LDR $0.0715$ $13.0605$ $2.77$ $6.86$ $88.38$ $0.0282$ $17.5467$ $3.15$ $6.68$ $97.21$
HDRCNN [12] $0.0246$ $23.6138$ $2.59$ $8.47$ $59.13$ $0.0060$ $30.6715$ $2.02$ $8.611$ $52.55$
MaskHDR [53] $0.0243$ $23.5503$ $2.52$ $8.54$ $58.33$ $0.0061$ $30.9510$ $1.86$ $8.62$ $54.83$
SingleHDR [36] $0.0178$ $24.0680$ $3.32$ $8.57$ $40.34$ $0.0092$ $23.6101$ $3.25$ $7.91$ $40.58$
ExpandNet [39] $0.0187$ $24.7158$ $2.10$ $8.48$ $58.11$ $0.0148$ $22.5799$ $4.56$ $8.06$ $56.46$
Exposure Diff [2] $0.0264$ $23.7384$ $8.23$ $7.88$ $47.94$ $0.0061$ $27.4963$ $2.51$ $8.38$ $63.51$
Ours $0.0131$ $25.3895$ $2.17$ $8.49$ $40.17$ $0.0041$ $28.7022$ $2.39$ $8.52$ $52.41$
Table 3: Quantitative evaluation of renderings lit by each method’s HDR image output. Results are color coded by best, second and third best. “P-PIQE” refers to PU21-PIQE and “H-VPD3” to HDR-VDP-3.
Relight HDR
Method MSE PSNR Ang. err H-VDP3 P-PIQE
LDR $0.049$ $14.04$ $1.83$ $7.30$ $90.00$
GlowGAN [59] $0.012$ $23.82$ $1.09$ $5.67$ $56.82$
Ours $0.013$ $25.05$ $1.14$ $8.49$ $37.56$
Table 4: Quantitative comparison to GlowGAN [59], trained on the “landscape” category, on selected landscape images from the Si-HDR dataset. Note that, unlike GlowGAN, our network is not category-specific. Results are color-coded by best, second and third best. “P-PIQE” refers to PU21-PIQE and “H-VDP3” to HDR-VDP-3.
We evaluate our novel light intensity estimation method (sec. 3.1) on the task of predicting bright light sources. We first show qualitative results on the filtered Si-HDR dataset in fig. 4. The figures illustrate both the luminance of the generated HDR maps, as well as a render of a sphere on a plane lit by the color HDR map. Since these images are not full 360° environment maps, we project them onto a partial hemisphere of 90° field of view, and use this as light source for rendering. The sphere rests on top of a plane positioned at the origin, facing the camera. The glossy sphere has an albedo of 1 (pure white) and a roughness coefficient of 0. We ensure that all HDR images used for lighting are exposed the same way by aligning the pixel averages with intensity between $0.2$ and $0.8$ with the ground truth. Our method’s predictions closely match the ground truth, particularly with bright light sources that cast hard shadows and create specularities. Please, see the supplementary material for qualitative results on the BtP-HDR dataset.
To further assess the accuracy of our method, we compute quantitative image comparison metrics on both renders (MSE, PSNR, and angular error) and the HDR images (HDR-VDP-3 [38] and PU21-PIQE [37]), compared to the ground truth in tab. 3. We observe that our method significantly outperforms existing methods in terms of both MSE and PSNR on the filtered Si-HDR dataset. On the BtP-HDR dataset, our method provides the best MSE score while offering a competitive PSNR value. We attribute this lower score to the fact that BtP-HDR was captured with a fixed white balance, leading to out-of-distribution color renditions. Our method predicts convincing content in saturated regions, which can lead to potentially different colors from the ground truth, as indicated by the angular error scores. Although our method is primarily designed for relighting, it also performs well in direct metrics.
To compare with GlowGAN [59], a subset of 23 landscape images from the Si-HDR dataset is selected, estimating HDR with the specialized checkpoints. As shown in tab. 4, our method is competitive even against the specialized GlowGAN.
@@
Chess room Small class
Relight HDR Relight HDR
Scene MSE PSNR P-PSNR MSE PSNR P-PSNR
PanoHDR-NeRF $0.0134$ $18.97$ $27.61$ $0.0077$ $21.68$ $28.57$
Ours $0.0026$ $30.50$ $30.29$ $0.0049$ $23.20$ $30.77$
Table 5: Quantitative evaluation of renderings lit by novel views of PanoHDR-NeRF [19] and Ours, as compared to the captured HDR ground truth panoramas provided in [19]. Our method achieves the best scores on both scenes for all metrics. “P-PSNR” refers to PU21-PSNR.
@@
4.4 Ablation
HDR predictions consistency
During HDR extrapolation, we process each image independently, which can lead to inconsistent predictions for the same light source. However, we found that the low frequencies of the spherical harmonics act as a form of regularization, ensuring consistency. As shown in fig. 5, even in extreme cases where HDR values were manually clamped to $1$, multi-view consistency is maintained after reconstructing the GS scene.
HDR
Refer to caption Refer to caption Refer to caption Refer to caption
Clamp
Refer to caption Refer to caption Refer to caption Refer to caption
3DGS
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 5: HDR predictions (top row) are manually clamped to simulate inconsistent HDR predictions (middle row). The spherical harmonics RGB representation in the GS reconstruction (bottom row) smooths the values and improves multi-view consistency. Images are displayed in linear EV-3 for clarity.
Using a different HDR estimation approach
To validate our Gaussian Splatting approach and HDR estimation method compared to PanoHDR-NeRF [19], we replace our light intensity estimation network with LaNet [67]. Tab. 6 shows a consistent decrease in quality with this change.
Relight HDR
Scene MSE PSNR PU21-PSNR
PanoHDR-NeRF $0.0134$ $18.97$ $27.61$
Ours+LaNet $0.0055$ $23.00$ $28.06$
Ours $0.0026$ $30.50$ $30.29$
Table 6: Quantitative comparaison between PanoHDR-NeRF, our method with LaNet [67] as HDR estimator, and our full method.
5 Applications
5.1 Virtual object insertion
The 3DGS representation allows for the direct composition of virtual objects, as shown in LABEL:fig:teaser and 6. We achieve these results by importing our 3DGS into Blender [9] with the Cycles renderer. We set Gaussians with intensity >1>\!1 as light emitters, and the rest as diffuse surfaces. This yields images with cast shadows and near-field lighting effects.
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 6: Examples of object insertions using GaSLight
5.2 Text-to-HDR environment
Instead of reconstructing the 3DGS scene on captured photographs, our method can be combined with recent generative approaches, which produce a 3DGS scene solely from text input. For example, we combine GaSLight with LucidDreamer [8], creating a spatially-varying HDR lighting representation from text, as shown in fig. 7.
Refer to caption Refer to caption Refer to caption
Prompt: “A gas lamp standing on a table in the center of a dark living room”
Refer to caption Refer to caption Refer to caption
Prompt: “A library with a window on one side”
Figure 7: Our method can easily be combined with recent prompt-based multi-view generative methods. Here we extend the text-to-scene LucidDreamer [8] method to GaSLight, resulting in text-to-lighting. 3D virtual objects (top: ducky, bottom: chair) are inserted to show the impact of the estimated HDR lighting.
@@ 6 Conclusion
Despite encouraging results, GaSLight suffers from some limitations. First, each Gaussian in the 3DGS representation is effectively considered to be a light source (e.g., when rendering, see sec. 5). This hinders the simulation of more complex lighting effects, such as interreflections or scattering. Here, it is likely that new raytracing-based approaches [42] could help. Second, our HDR light intensity estimation network is not trained to achieve multi-view consistency: it is therefore possible that different content is extrapolated for two nearby views. While the spherical harmonics in the 3DGS representation alleviate this somewhat and can realistically blend inconsistent predictions, future work could enforce such consistency, for example, through score distillation sampling [48]. We also note that, unlike most lighting estimation methods, which predict light sources outside the field of view of the image, we only recover the lighting for visible parts of the scene. Combining our method with view extrapolation approaches (e.g., [8, 45]) is a promising future direction. Additionally, although our method can produce HDR environment maps for fast rendering, accurate near-field lighting effects require ray tracing due to the high density of Gaussians generated, which is impractical for real-time use. Finally, our current implementation, which relies on InstantSplat [14], does not scale well to large scenes. This is somewhat orthogonal to our goal of capturing lighting—we are hopeful that new developments (e.g., [31]) could help alleviate this while being compatible with our method.
We present GaSLight, a novel Gaussian Splatting-based HDR Lighting representation. Our method employs a diffusion-based LDR-to-HDR model that enhances the dynamic range of LDR images to an unsaturated HDR, even in the presence of very bright light sources. We then fit a 3DGS model on the predicted HDR images, resulting in a lighting representation producing spatially-varying effects, high-frequency reflections, and hard cast shadows, which can be easily integrated into existing rendering engines.
Acknowledgements This work was supported by Sentinel North, NSERC grants RGPIN 2020-04799 and ALLRP 557208-20, and the Digital Research Alliance Canada.